# VLsI: Verbalized Layers-to-Interactions from Large to Small Vision Language Models

### v1

@@ Abstract
VLsI across ten challenging vision-language benchmarks, achieving notable performance gains (11.0% for 2B and 17.4% for 7B) over GPT-4V without the need for model scaling, merging, or architectural changes. Project page is now accessible.
@@ 1 Introduction
VLsI introduces a layer-wise approach where each intermediate layer generates verbal responses in natural language space, enhancing interpretability and alignment with larger models. This is achieved through a three-step process: (1) the verbalization step, which uses “verbalizers” to project intermediate features into the language space, making them interpretable as text-based responses; (2) the interaction step, which performs adaptive layer matching to align the reasoning progression between large and small VLMs; and (3) the reinforcement step, which finetunes the distilled VLMs for task-specific instruction-following responsiveness.
@@
VLsI’s effectiveness across ten challenging benchmarks, demonstrating significant performance gains of 11.0% (2B model) and 17.4% (7B model) over GPT-4V. Notably, these improvements are achieved without increasing model size, merging modules, or modifying the architecture. Our contributions are as follows:
VLsI, showing (a) the verbalization step and (b) the interaction step. (a) In the verbalization step, intermediate layers in both the large- and small-backbone VLMs are equipped with a “verbalizer”, allowing their outputs to be projected into natural language space. Autoregressive loss is applied to align these verbalized outputs with the target responses. (b) In the interaction step, each intermediate layer in the small-backbone VLM searches for a matching layer in the large backbone VLM within a specified range. For example, once the 2 layer of the small VLM is matched with the 4 layer in the large VLM, the next matching search for the 3 layer in the small VLM will proceed from the 5 to the 7 layers of the large VLM, ensuring progressive alignment.
@@
VLsI achieves high performance without adding extra modules or architectural modifications, making it a practical and deployable solution for on-device applications in resource-constrained environments.
VLsI is easy to implement and adaptable across different model architectures, showing significant gains not only with Qwen2-VL but also with LLaVA-OV, where it achieves a 19.7% improvement in 2B and 7B model sizes (Qwen2-VL), and a 34.5% improvement in 0.5B and 7B model sizes (LLaVA-OV) on challenging benchmarks like MMB, MM-Vet, and MMMU.
This work not only advances the state of VLMs in a computationally efficient manner but also sets a precedent for deploying high-performing, low-complexity models in real-world applications, bringing us a step closer to scalable, practical artificial intelligence.
@@ 2 Related Works
The emergence of visual instruction tuning: LLaVA [60] and InstructBLIP [17] initially brings in not only introducing slight variations of VLMs [44, 7, 3, 106, 50, 16, 101], but also curating high-quality visual instruction samples [8, 6, 93, 30, 86, 103, 26]. Since that time, there has been a growing interest in enhancing visual understanding; thus the simple visual input technique of enlarging image resolution or dividing images into smaller sections with fixed or dynamic rules has got standardized [61, 69, 56, 89, 11, 31]. Furthermore, merging additional visual encoders [25, 75, 43, 99] or multiple computer vision models [12, 71, 91, 21] into LLMs have also become a major focus [65, 32, 107, 48, 49, 77]. Besides, Meteor [47] employs an additional rational projector that embeds multifaceted reasoning information to cover diverse capabilities, including chart, diagram, document, and math reasoning. More recently, Cambrian-1 [84], LLaVA-OneVision [52] InternVL2 [11], Molmo [19], and Qwen2-VL [87] have released large scale models in order to follow or surpass the performances of closed-source VLMs.
While these advancements are both rapid and impactful, relatively few studies focus on achieving high-performing yet efficient-scale VLMs within limited architectures. Furthermore, many current approaches rely heavily on GPT-based instruction datasets, and only the final layer learns the target responses in visual instruction tuning. On the other hand,
@@
In the field of lightweight LLMs, MobiLlama [83], OpenELM [70], MobileLLM [63] have leveraged various engineering techniques such as shared feed-forward network (FFN) design, layer-wise scaling, and embedding-language head sharing to efficiently reduce model parameters. Their primary objective is not to close the gap with closed-source LLMs but rather to reduce parameters impacting less performance degradation. In the field of VLMs, there are variations about how to utilize pretrained lightweight LLMs [102, 2, 33] in order to make efficient-scale VLMs [105, 13, 14, 57, 76, 104], but these works are also not a fundamental solution to embed more vision-language knowledge within limited structures. Notably, two VLM approaches, TroL [46] and Phantom [45], aim to expand learning own capabilities within limited structures by doubling forward propagation steps and enlarging the latent dimension without physically increasing model sizes, thereby showing large improvements. Unfortunately, these approaches face limitations such as key-value cache storage constraints and extensive architectural modifications, which may hinder direct application to real-world scenarios. Besides, a few works in distilling LLMs [72, 80] have emerged, but they use in the end for the final layer distillation. For VLMs, LLaVA-MoD [78] and LLaVA-KD [4] have also been recently proposed, but they use the same way.
On the other hand,
VLsI makes layer-wise distillation process where we leverage natural language in order to make small VLMs mimic the reasoning progression of large VLMs across layers. We hope that incorporating natural language will facilitate smoother communication between large and small VLMs, alleviating the complexities of direct feature alignment.
@@ 3 VLsI: Verbalized Layers-to-Interactions
VLsI comprises two main components: the backbone VLM and verbalizers. For the backbone VLM, we use Qwen2-VL [87], selected for its high performance on the OpenVLM-Leaderboard [15]. The verbalizer consists of a simple FFN [85] and the language head from the backbone VLM. To distinguish between FFN components, we refer to the FFN within the verbalizer as “verb-FFN” to avoid confusion with the FFN in the LLM transformer decoder block. This design is inspired by the recent speculative decoding paradigm [55], which demonstrates the effectiveness of using a smaller LLM constructed with a frozen word embedding and the language head of larger LLMs to emulate the performance of those larger models. Building on this insight, we incorporate the language head of the backbone VLM into its intermediate layers. Specifically, the verb-FFN is placed between each intermediate layer and the language head, introducing trainable parameters that allow effective projection into the language space. In other words, the verb-FFN enables outputs from the intermediate layers to project to the language space via the language head. In the following sections, we detail the three critical training stages: verbalization, interaction, and reinforcement.
@@ 3.2 Interaction Step
3.3 Reinforcement Step
In the final stage of our distillation framework, we finetune the entire small-backbone VLM, including word embeddings, multi-head attention, FFN, and the language head on the visual instruction dataset in a supervised-learning manner. This reinforcement step is inspired by pruning-based distillation methods [72, 80], which require additional training after pruning to counteract potential performance drops from structural changes. While our interaction step does not alter the model’s structure, fully absorbing the rich information from the large VLM remains challenging for the small VLM within a single interaction step. The reinforcement step allows the small VLM to better align and integrate the acquired knowledge across layers, enhancing response quality through autoregressive loss. We avoid optimizing autoregressive loss during the interaction step (where KL divergence loss is optimized instead) as this can degrade performance, as demonstrated in our ablation study (Sec. 4). Therefore, to improve task-specific instruction-following responsiveness and accuracy, we incorporate the reinforcement step with autoregressive loss to further finetune the small-backbone VLM.
@@ 4.1 Implementation Details
| VLMs | RS | MMB | BLINK | MM-Vet | MMMU |
@@ 4.2 Verbalization on
VLsI’s layer-wise verbalization, where each stage of verbal responses helps the small-backbone VLM align with the larger one. Additional examples of
VLsI’s verbalization are available in Appendix B, highlighting its capacity to interpret layer-wise verbal responses.
@@ 4.3 Comparison on Evaluation Benchmarks
VLsI achieves higher performance while maintaining an efficient model scale. Furthermore, Sec. 4.1 compares
VLsI’s performance with (a) open-source VLMs that incorporate multiple vision encoders, computer vision models, and additional rationale projector; (b) VLMs with modified architectures [46, 45] and various larger open- and proprietary closed-source VLMs. To evaluate where the effectiveness comes from, we analyze six key factors detailed in Fig. 4 and Tab. 4. Appendix C explains the detailed settings for ablation studies. These results indicate that (1) utilizing a more capable large-backbone VLM provides substantial performance benefits, suggesting that the choice of backbone significantly impacts the transfer of knowledge; (2) using larger-backbone VLM gets benefits; (3) KL divergence is more effective during interaction step than cross-entropy, and simultaneously using last-layer distillation boosts performances; and (4) the reinforcement step is crucial for further performance gains, consistent with findings in [72, 80]. Appendix D represents
@@ Appendix A Description of Numeorus Evaluation Benchmarks
Appendix B VLsI’s Verbalization Examples
Appendix C Comprehensive Experimental Setup for Ablation Studies
@@
Tab. 4(b) and Tab. 4(c)
collectively illustrate the impact of incorporating the reinforcement step (RS). Table (b) highlights the significant performance gains achieved by applying RS, while Table (c) examines whether these improvements result from the fine-tuning effects of 2.9 million visual instruction tuning samples. The results demonstrate that
VLsI’s performance becomes markedly superior to Qwen2-VL as RS training progresses from 50% to 100%, confirming that RS plays a much more critical role in driving performance enhancements than the contribution of the visual instruction tuning samples alone.
@@
VLsI’s performances. The results clearly demonstrate that KL divergence (KLD) operations applied to both intermediate and last layers yield the best performances, confirming the effectiveness of KLD over cross-entropy (CE) or L2 for intermediate layer alignment and final layer interaction. To further assess the effectiveness of verbalization in transferring the knowledge via distillation, we conduct an ablation study by removing the language head in verbalizer (vL-head). In this setup, the verbalization step is skipped, and the interaction step is first performed by aligning the hidden dimensions of the verb-FFNs in large- and small-backbone VLM. Here, the verb-FFN’s hidden dimension in large-backbone VLM is kept but the hidden dimension in small-backbone VLM’s verb-FFN is enlarged to match the large one. In addition, the verb-FFNs of large- and small-backbone VLM are interacted with only L2, which means that small-backbone VLM’s verb-FFNs try to naively follow those of large-backbone VLM. After interaction step, reinforcement step is equally conducted. However, as shown in the above table, this naive approach results in significantly lower performance compared to the version equipped with the language head. Interestingly, using only last layer distillation by L2 and KLD provides more benefits than using intermediate distillation without vL-Head. These findings suggest that directly imitating outputs from the large-backbone VLM, without verbalization from vL-Head, introduces instability and can lead to suboptimal results, highlighting the critical role of the language head in achieving effective distillation.
@@
highlights the effectiveness of various components in the matching strategy. Random Index and Bottom-1/3 Index, yield lower scores, underscoring the limitations of simpler selection mechanisms. Note that, the above algorithms represent the their detailed experimental setup. Multinomial sampling provides improvements and incorporating the Search Range further enhances performances (e.g., 83.5 on MMB and 69.8 on MM-Vet). Adding order preservation results in a significant leap, particularly on BLINK (59.2) and MM-Vet (75.2), demonstrating the importance of maintaining matched indices’ sequence alignment during interaction step. Finally, using all together with adaptive temperature achieves the best results across all benchmarks (e.g., 86.3 on MMB and 69.3 on MMMU), showcasing its ability to dynamically control the distribution. These exploration underscore the necessity of advanced sampling strategies and adaptive mechanisms for maximizing the efficiency of transferring the knowledge.
@@
Appendix D VLsI’s Text Generation Quality

### v2

@@ Abstract
VLsI across ten challenging vision-language benchmarks, achieving notable performance gains (11.0% for 2B and 17.4% for 7B) over GPT-4V without the need for model scaling, merging, or architectural changes. Project Page.
@@ 1 Introduction
VLsI introduces a layer-wise approach where each intermediate layer generates verbal responses in natural language space, enhancing interpretability and alignment with larger models. This is achieved through a three-step process: (1) the verbalization step, which uses “verbalizers” to project intermediate features into the language space, making them interpretable as text-based responses; (2) the interaction step, which performs adaptive layer matching to align the reasoning progression between large and small VLMs; and (3) the SFT step, which finetunes the distilled VLMs for task-specific instruction-following responsiveness.
@@
VLsI’s effectiveness across ten challenging benchmarks, demonstrating significant performance gains of 11.0% (2B model) and 17.4% (7B model) over GPT-4V. Notably, these improvements are achieved without increasing model size, merging modules, or modifying the architecture. Our contributions are as:
VLsI, showing (a) the verbalization step and (b) the interaction step. (a) In the verbalization step, intermediate layers in both the large- and small-backbone VLMs are equipped with a “verbalizer”, allowing their outputs to be projected into natural language space. Autoregressive loss is applied to align these verbalized outputs with the target responses. (b) In the interaction step, each intermediate layer in the small-backbone VLM searches for a matching layer in the large backbone VLM within a specified range. For example, once the 2 layer of the small VLM is matched with the 4 layer in the large VLM, the next matching search for the 3 layer in the small VLM will proceed from the 5 to the 7 layers of the large VLM, ensuring progressive alignment. Note that, “fireball” icon represents parameters that are actively trained or updated, while the “ice cube” icon indicates untrained parameters.
@@
We present sampling distribution-based dynamic layer matching before distillation, in order to address the layer number gap between large and small VLMs.
VLsI is easy to implement and adaptable across various model architectures, making it practical and deployable for real-world applications.
@@ 2 Related Works
The emergence of visual instruction tuning: LLaVA [65] and InstructBLIP [18] initially brings in not only introducing slight variations of VLMs [48, 8, 3, 112, 54, 17, 107], but also curating high-quality visual instruction samples [9, 7, 99, 32, 91, 109, 27]. Since that time, there has been a growing interest in enhancing visual understanding; thus the simple visual input technique of enlarging image resolution or dividing images into smaller sections with fixed or dynamic rules has got standardized [66, 74, 60, 94, 12, 33]. Furthermore, merging additional visual encoders [26, 80, 46, 105] or multiple computer vision models [13, 76, 97, 22] into LLMs have also become a major focus [70, 34, 113, 52, 53, 82]. Besides, Meteor [51] employs an additional rational projector that embeds multifaceted reasoning information to cover diverse capabilities, including chart, diagram, document, and math reasoning. More recently, Cambrian-1 [89], LLaVA-OneVision [56] InternVL2 [12], Molmo [20], and Qwen2-VL [92] have released large scale models in order to follow or surpass the performances of closed-source VLMs. While these advancements are both rapid and impactful, relatively few studies focus on achieving high-performing yet efficient-scale VLMs within limited architectures. Furthermore, many current approaches rely heavily on GPT-based instruction datasets, and only the final layer learns the target responses in visual instruction tuning. In contrast,
@@
In the field of lightweight LLMs, MobiLlama [88], OpenELM [75], MobileLLM [68] have leveraged various engineering techniques such as shared feed-forward network (FFN) design, layer-wise scaling, and embedding-language head sharing to efficiently reduce model parameters. Their primary objective is not to close the gap with closed-source LLMs but rather to reduce parameters impacting less performance degradation. In the field of VLMs, there are variations about how to utilize pretrained lightweight LLMs [108, 2, 35] in order to make efficient-scale VLMs [111, 14, 15, 62, 81, 110], but these works are also not a fundamental solution to embed more vision-language knowledge within limited structures. Notably, two VLM approaches, TroL [50] and Phantom [49], aim to expand learning own capabilities within limited structures by doubling forward propagation steps and enlarging the latent dimension without physically increasing model sizes, thereby showing large improvements. Unfortunately, these approaches face limitations such as key-value cache storage constraints and extensive architectural modifications, which may hinder direct application to real-world scenarios. Besides, a few works in distilling language models [61, 77, 85] have emerged, and TED [61] utilizes language head but it is not a version of LLM covering general tasks, and it does not handle layer number gap between large and small models. The other studies simply use the final layer distillation. For VLMs, LLaVA-MoD [83] and LLaVA-KD [4] have also been recently proposed, but they use the same way. Align-KD [28] distills both vision features and logits from the language head, while MoVE-KD [6] employs multiple vision encoders for large VLMs and a mixture of LoRA [40] for small VLMs. Regarding the distillation metric, DistiLLM [47] and MiniLLM [36] emphasize the effectiveness of reverse KL divergence, whereas LLAVADI [95] and Minitron [77] demonstrate that normal KL divergence is more effective. On the other hand,
VLsI makes layer-wise distillation process where we leverage natural language in order to make small VLMs mimic the reasoning progression of large VLMs across layers. We hope that incorporating natural language will facilitate smoother communication between large and small VLMs, alleviating the complexities of feature alignment.
@@ 3 VLsI: Verbalized Layers-to-Interactions
VLsI comprises two main components: the backbone VLM and verbalizers. For the backbone VLM, we use Qwen2-VL [92], selected for its high performance on the OpenVLM-Leaderboard [16]. The verbalizer consists of a simple FFN [90] and the language head from the backbone VLM. To distinguish between FFN components, we refer to the FFN within the verbalizer as “verb-FFN” to avoid confusion with the FFN in the LLM transformer decoder block. This design is inspired by the recent speculative decoding paradigm [59], which demonstrates the effectiveness of using a smaller LLM constructed with a frozen word embedding and the language head of larger LLMs to emulate the performance of those larger models. Building on this insight, we incorporate the language head of the backbone VLM into its intermediate layers. Specifically, the verb-FFN is placed between each intermediate layer and the language head, introducing trainable parameters that allow effective projection into the language space. In other words, the verb-FFN enables outputs from the intermediate layers to project to the language space via the language head. In the following sections, we detail the three critical training stages: verbalization, interaction, and SFT.
@@ 3.2 Interaction Step
3.3 Supervised Finetuning (SFT) Step
In the final stage of our distillation framework, we finetune the entire small-backbone VLM, including word embeddings, multi-head attention, FFN, and the language head on the visual instruction dataset in a supervised-learning manner. This SFT step is inspired by pruning-based distillation methods [77, 85], which require additional training after pruning to counteract potential performance drops from structural changes. While our interaction step does not alter the model’s structure, fully absorbing the rich information from the large VLM remains challenging for the small VLM within a single interaction step. The SFT step allows the small VLM to better align and integrate the acquired knowledge across layers, enhancing response quality through autoregressive loss. We avoid optimizing autoregressive loss during the interaction step (where KL divergence loss is optimized instead) as this can degrade performance, as demonstrated in our ablation study (Sec. 4). Therefore, to improve task-specific instruction-following responsiveness and accuracy, we incorporate the SFT step with autoregressive loss to further finetune the small-backbone VLM.
@@ 4.1 Implementation Details
| VLMs | SFT | MMB | BLINK | MM-Vet | MMMU |
@@ 4.2 Verbalization on
VLsI’s layer-wise verbalization, where each stage of verbal responses helps the small-backbone VLM align with the larger one. Note that, Appendix B shows the loss changes during verbalization. Additional examples of
VLsI’s verbalization are available in Appendix C, highlighting its capacity to interpret layer-wise verbal responses.
@@ 4.3 Comparison on Evaluation Benchmarks
VLsI achieves higher performance while maintaining an efficient model scale. Furthermore, Tab. 3 compares
VLsI’s performance with (a) open-source VLMs that incorporate multiple vision encoders, computer vision models, and additional rationale projector; (b) VLMs with modified architectures [50, 49] and various larger open- and proprietary closed-source VLMs. To evaluate where the effectiveness comes from, we analyze six key factors detailed in Fig. 4 and Tab. 4. Appendix D explains the detailed settings for ablation studies. These results indicate that (1) utilizing a more capable large-backbone VLM provides substantial performance benefits, suggesting that the choice of backbone significantly impacts the transfer of knowledge; (2) using larger-backbone VLM gets benefits; (3) KL divergence is more effective during interaction step than cross-entropy, and simultaneously using last-layer distillation boosts performances; and (4) additional SFT step is crucial for further performance gains, consistent with findings in [77, 85]. Appendix E represents
@@ Appendix A Description of Numeorus Evaluation Benchmarks
Appendix B Verbalization Loss Changes
| Target Layer Index | 10% | 30% | 50% | 70% | 90% | 100% |
| #1 | 14.28 | 11.73 | 9.87 | 6.78 | 3.54 | 3.30 |
| #2 | 12.60 | 10.48 | 8.04 | 5.98 | 3.10 | 2.90 |
| #3 | 11.22 | 9.32 | 7.02 | 5.05 | 2.63 | 2.37 |
| #4 | 10.48 | 8.11 | 6.41 | 4.57 | 2.35 | 2.18 |
| #5 | 9.34 | 7.55 | 5.81 | 3.54 | 1.82 | 1.79 |
| #6 | 8.16 | 6.18 | 4.27 | 2.09 | 1.56 | 1.47 |
| #7 | 7.39 | 3.85 | 2.02 | 1.03 | 0.85 | 0.79 |
Appendix C VLsI’s Verbalization Examples
Appendix D Comprehensive Experimental Setup for Ablation Studies
@@
Tab. 4(b)
collectively illustrate the impact of incorporating the SFT step. Table (b) highlights the significant performance gains achieved by applying additional SFT. Interestingly, the performances without additional SFT step are not satisfactory when doing greedy decoding, but once we utilize beam decoding (=5), its performances are dramatically enhanced. This suggests that the interaction step implicitly alters and expands the probability space over the language head, while SFT aligns this expanded space to better fit instruction-following. Thus, significant improvements arise from the combination of interaction and SFT step.
Tab. 4(c)
It examines whether these improvements result from the fine-tuning effects of 2.9 million visual instruction tuning samples. The results demonstrate that
VLsI’s performance becomes markedly superior to Qwen2-VL as SFT training progresses from 50% to 100%, confirming that additional SFT plays a much more critical role in driving performance enhancements than the contribution of the visual instruction tuning samples alone.
@@
VLsI’s performances. The results clearly demonstrate that KL divergence (KLD) operations applied to both intermediate and last layers yield the best performances, confirming the effectiveness of KLD over cross-entropy (CE) or L2 for intermediate layer alignment and final layer interaction. To further assess the effectiveness of verbalization in transferring the knowledge via distillation, we conduct an ablation study by removing the language head in verbalizer (vL-head). In this setup, the verbalization step is skipped, and the interaction step is first performed by aligning the hidden dimensions of the verb-FFNs in large- and small-backbone VLM. Here, the verb-FFN’s hidden dimension in large-backbone VLM is kept but the hidden dimension in small-backbone VLM’s verb-FFN is enlarged to match the large one. In addition, the verb-FFNs of large- and small-backbone VLM are interacted with only L2, which means that small-backbone VLM’s verb-FFNs try to naively follow those of large-backbone VLM. After interaction step, SFT step is equally conducted. However, as shown in the above table, this naive approach results in significantly lower performance compared to the version equipped with the language head. Interestingly, using only last layer distillation by L2 and KLD provides more benefits than using intermediate distillation without vL-Head. These findings suggest that directly imitating outputs from the large-backbone VLM, without verbalization from vL-Head, introduces instability and can lead to suboptimal results, highlighting the critical role of the language head in achieving effective distillation.
@@
highlights the effectiveness of various components in the matching strategy. Random Index, Uniform Index, and Bottom-1 or 3 Index, yield lower scores, underscoring the limitations of simpler selection mechanisms. Note that, the above and below algorithms represent the their detailed experimental setup. Multinomial sampling provides improvements and incorporating the Search Range further enhances performances (e.g., 83.5 on MMB and 69.8 on MM-Vet). Adding order preservation results in a significant leap, particularly on BLINK (59.2) and MM-Vet (75.2), demonstrating the importance of maintaining matched indices’ sequence alignment during interaction step. Finally, using all together with adaptive temperature achieves the best results across all benchmarks (e.g., 86.3 on MMB and 69.3 on MMMU), showcasing its ability to dynamically control the distribution. These exploration underscore the necessity of advanced sampling strategies and adaptive mechanisms for maximizing the efficiency of transferring the knowledge.
@@
Appendix E VLsI’s Text Generation Quality